OpenStudy (anonymous):

The top of a tree makes angles s and t with Points K and L on the ground, respectively, such that the angles are complementary. Point K is x meters and Point L is y meters from the base of the tree. In terms of x and y, find the height of the tree. Include your work.

OpenStudy (dumbcow):

 use tangent ratio: \[\tan (s) = \frac{x}{h}\] \[\tan(s) = \frac{h}{y}\] set equal and solve for h \[\frac{x}{h} = \frac{h}{y} \rightarrow h^2 = xy\] \[h = \sqrt{xy}\]